40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed May 15 19:00:26 2019, anchal, DailyProgress, South Cavity, Changes on South Path towards increasing power on FSS RFPD  
    Reply  Thu May 16 16:34:30 2019, anchal, DailyProgress, North Cavity, Changes on North Path towards increasing power on FSS RFPD  
Message ID: 2346     Entry time: Wed May 15 19:00:26 2019     Reply to this: 2348
Author: anchal 
Type: DailyProgress 
Category: South Cavity 
Subject: Changes on South Path towards increasing power on FSS RFPD  

For gaining about 3.45 mW of light after reflection from cavity on resonance when mode matching is about 60%, 8.63 mW light needs to fall at the cavity. There were three major obstacles in doing this which were resolved as follow:

  • PMC PDH becoming unstable: Too much light at the reflection of PMC PDH was making the locks unstable. This would be primarily due to the railing of RF amplifier in the RFPDs. To circumvent this problem, as Rana suggested, I discarded about 60% of light reflecting from South PMC by placing a beamsplitter (BS1-1064-50-1025-45S) instead of the last mirror that directs light towards it. Since it is ideal for 's' polarized light while the reflected light from PMC is 'p' polarized, the beam splitter is discarding more than 50% which is desirable. This allowed me to increase the light intensity much more than before.
  • PMC PDH Modulation Depth:  I found that the modulation depth for PMC PDH was really high. I measured it to be ~0.63 by directly measuring the transmitted light level of carrier and sidebands. I wasn't doing this before and the modulation actuation slope was too uncertain in the datasheet, hence I messed this up. Now, I have inserted a 9 dB attenuator to the modulation signal that goes to the resonant EOM for PMC PDH and now the modulation depth is ~0.23.
  • Decreasing Cavity Relf RFPD DC Transimpedance: SN010, South Cavity Refl RFPD had a DC transimpedance of 2222 Ohm which was getting saturated on 6.54 mW of light on the detector. This was making it hard for the autolocker scripts which rely on DC reflection value to estimate lock quality. Hence, I reduced this DC transimpedance to 1022 by changing R11 to 22 Ohms. Now we have more lee-way.

Also, I have finally started an ATF wiki page to keep documentation of all RFPDs and changes to it. I'll make sure this page is updated and this page should be the reference point in future.

https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:rfpd


On another note, I'm seeing a steady decline of mode matching on both paths even when I'm staying far away from the alignment. Something is going on which needs to be fixed. After a discussion with awade, following are the suspects:

  • Low nitrogen level in table legs: I have checked this already and it is between 45-50 psi, so we are good. I'm writing this down mostly for note taking purpose.
  • Cavity height changes: Currently, no active temperature control is in place. I'm seeing cavity resonance varying widely from one day to another. So the clamp of the cavities might be expanding/contracting over days changing the alignment into the cavity. Fixing vacuum can temperature control is important definitely.
  • Periscope or other steering mirrors relaxing: One reason might be that the periscope mirrors or other steering mirrors between Faraday isolator and cavity are slowly relaxing. I doubt this is the case because then it should have stopped after a certain time, but I'm seeing a steady decline in mode matching for 3 days now.

I'll fix the mode matching again before leaving today and will see if any big changes happen overnight.

ELOG V3.1.3-